### v2

Abstract
Singularly perturbed dynamical systems play a crucial role in climate dynamics and plasma physics. A powerful and well-known tool to address these systems is the Fenichel normal form, which significantly simplifies fast dynamics near slow manifolds through a transformation. However, this normal form is difficult to realize in conventional numerical algorithms. In this work, we explore an alternative way of realizing it through structure-preserving machine learning. Specifically, a fast-slow neural network (FSNN) is proposed for learning data-driven models of singularly perturbed dynamical systems with dissipative fast timescale dynamics. Our method enforces the existence of a trainable, attracting invariant slow manifold as a hard constraint. Closed-form representation of the slow manifold enables efficient integration on the slow time scale and significantly improves prediction accuracy beyond the training data. We demonstrate the FSNN on examples including the Grad moment system, two-scale Lorenz96 equations, and Abraham-Lorentz dynamics modeling radiation reaction of electrons.
keywords:
Singularly perturbed dynamical systems, slow manifold, model reduction, closures
1 Introduction
In this article, a new architecture called the Fast-Slow Neural Network (FSNN) is developed for learning data-driven models of singularly perturbed dynamical systems with dissipative fast timescale dynamics. We refer to our approach as an intelligent attractor since the method enforces the existence of a trainable, attracting invariant slow manifold [1] as a hard constraint. The constraint is imposed by way of a novel neural network architecture with structured weight matrices. By leveraging the neural ODE framework from [2], our technique produces deterministic evolution equations that approximately reproduce trajectory training data. When the training data comes from an unknown dynamical system, the method achieves data-driven full-order model discovery with stability guarantees due to provable existence of the slow manifold. In short, our architecture is, by design, a singularly perturbed dynamical system which contains an attractor.
Various dynamical systems encountered in nature exhibit a timescale separation such that the fast timescale is dissipative. Prominent examples include the kinetic dynamics of strongly-collisional gasses and plasmas, and radiative transfer in optically-thick media. For ordinary differential equations (ODEs) of this type, Fenichel theory [3] reveals that fast timescale dissipation implies existence of an invariant slow manifold that attracts nearby trajectories and governs the long-term dynamics. For partial differential equations (PDEs), or more generally integro-differential equations, Fenichel theory generally only applies formally, but its predictions are often valid anyway. For PDEs, inertial manifold theory [4] has been studied and shown to be useful in applications such as model reductions of fluid dynamics.
Nonlinear dimensionality reduction inspired by inertial manifold theory has a rich history (see [5] and the references therein) originating in the 1980s and 1990s [6, 4, 7, 8, 9, 10, 11, 12]. Data-driven approaches for discovering inertial manifolds in dynamical systems and using separations of scales to evolve a lower-dimensional reduced order model first materialized in the 90s [13, 14, 15, 16, 17, 18, 19] and recently reemerged in popularity due to boom in machine-learning-based modeling [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. The work in this manuscript tackles a special class of systems with attracting invariant sets, namely fast-slow systems with dissipative fast timescale dynamics [1]. Unlike previous data-driven approaches, the FSNN is a full-order model which, using neural network architectures with special properties, strictly enforces the existence of an attracting slow manifold. This method may be viewed as a major step along a pathway in developing data-driven model discovery techniques with embedded invariant manifolds. In future work, we plan to address the general case where explicit timescale separation is not used as a crutch.
The secondary goal of the proposed network is to discover an efficient and asymptotic-preserving reduced-order model. In the common methodology of physics-based model reduction, one first decides a latent space (e.g. lower-order moments of a distribution) and then looks for a closure using tools such as asymptotic analysis. Many complicated closures have been discovered through this approach. The classical examples in rarefied gases and plasma physics include Grad’s moment method [33], Braginskii closure [34] and Hammett-Perkins closure [35]. The challenge of this methodology is that an exact closure may not exist. The accuracy of the approximate closure is related to the dimension of the latent space. In contrast, Fenichel theory guarantees the existence of an exact closure in the neighborhood of the singular limit of the equations in the form of a diffeomorphic coordinate transformation. This transformation exposes an invariant manifold by flattening the fast variable space. Often in practice, as asymptotic expansion is the best analytical approximation to the exact closure. In our work, we leverage structured neural network architectures and optimization to learn an approximation to the exact closure that is hidden in the dynamics. The data-driven approach described in this work can discover such a transformation only using short-term trajectory data. Therefore, this approach may be useful when the underlying dynamical system is known but the targeting latent space or a formal closure are not known.
The design of the FSNN architecture is motivated by the first few theorems discovered by Fenichel [36]. The components of the architecture are constructed using structured weight matrices. The coordinate transformation into Fenichel normal form is based on an invertible coupling flow network [37] in which we introduced the additional structure-preserving property of bi-Lipschitz control through the use of bi-Lipschitz affine transformation (bLAT) layers. This network and its inverse can be evaluated using closed-form expressions with identical computational complexity. We note that “vanilla” neural networks are generally non-invertible [38, 39, 40]. We also introduce the Schur form network, which is a novel architecture based on the Schur decomposition for controlling the locations of eigenvalues in the complex plane, to parameterize the linear term in the dynamics. Additionally, a low-rank bilinear map network is included to increase the approximation power in the fast variable space. Our approach is then integrated using adaptive time stepping that treats the linear term implicitly, further leveraging back-substitution on the Schur form. Some of the structures introduced are related to the idea of trivializing manifold optimization through parameterizations [41]. Structure-preserving flow maps through machine learning are an active area of research, and include such as the symplectic neural networks for Hamiltonian dynamics [42, 43]. Previous studies on flow map learning for multi-scale dynamical systems include [44] and [45]. In particular, [45] simultaneously preserves symplecticity and adiabatic invariance in a multi-scale setting.
Before going into a detailed presentation of our new method, it is worth highlighting an important shortcoming it suffers from, and how that shortcoming might be mitigated in practice. The method assumes that the dimension of the attracting invariant manifold is known in advance. While this is not a limitation in the surrogate modeling context, where the (expensive to simulate) dynamical system is known in detail, it does present a moderate challenge in the model discovery context. As such, when performing data-driven model discovery using our method, it will be necessary to first determine the dimension of the slow manifold, either through physical reasoning, or through application of auxiliary data mining techniques such as diffusion maps [46], autoencoders [26], or other techniques from reduced order modeling [5]. For a review on various manifold learning techniques, we refer the reader to [47]. In many cases, including the examples considered in this manuscript, the form of the slow manifold and therefore its dimensionality are known in the singular limit $\epsilon=0$ as a power series expansion of $\epsilon$. However, the exact form for $\epsilon>0$ and the transformation into Fenichel normal form is unknown. In order to learn the slow manifold effectively, a sufficient number of trajectories for various choices of $\epsilon$ and initial conditions are necessary. In the final example of the paper, we apply our network in the nontrivial setting of learning the slow manifold for the Abraham-Lorentz equations which, for trajectories starting off of the slow manifold, are unstable forward in time.
The remainder of the manuscript is organized as follows. A review of fast-slow dynamical systems is presented in Section 2. The fast-slow neural network architecture is described in detail in Section 3. The numerical implementation of the architecture is presented in Section 4. Results from applying the architecture to various examples are shown in Section 5. Concluding remarks are made in Section 6.
2 Fast-Slow Dynamics and the Fenichel Normal Form
In this paper, we develop a novel machine learning (ML)-inspired architecture, the Fast-Slow Neural Network (FSNN). FSNNs are a parameterization for a class of dynamical systems exhibiting multiscale behavior in time that, to leading order, are dissipative on the fast time scale. Fast-slow systems, which exhibit both short-term and long-term behavior, are relevant to many physical systems observed in nature, including the climate and plasma systems. In this section, we motivate the FSNN architecture with a review of the theory of fast-slow systems.
[…]
A fast-slow system is an ODE of the form [36, 1]
(1a) $\epsilon\frac{d}{d\tau}y$ $=f(x,y,\epsilon),$
(1b) $\frac{d}{d\tau}x$ $=g(x,y,\epsilon),$
where $x=x(\tau)\in\mathbb{R}^{N_{x}}$, $y=y(\tau)\in\mathbb{R}^{N_{y}}$, $f$ and $g$ are smooth functions of their arguments, and $D_{y}f(x,y,0)$ is invertible for all $(x,y)$ where $f(x,y,0)=0$.
In the above definition, $y$ and $x$ are the fast and slow variables, respectively, over corresponding timescale variables $t$ and $\tau=\epsilon t$.
In some instances, a dynamical system may exhibit a fast-slow split, but a representation for the fast and slow variables is unknown. Consider the following representation of a dynamical system for $z(\tau)\in\mathbb{R}^{N}$, $\epsilon>0$,
(2) $\epsilon\frac{d}{d\tau}z$ $=U(z,\epsilon),$
where $U$ is a sufficiently smooth function of $z$ and $\epsilon$.
[…]
The dynamical system (2) exhibits a fast-slow split if there is an $\epsilon$-dependent invertible transformation of the dependent variables into fast-slow coordinates $(x,y)$, $x\in\mathbb{R}^{N_{x}}$, $y\in\mathbb{R}^{N_{y}}$, $N=N_{x}+N_{y}$, where the dynamics are described by a fast-slow system.
In this paper, we restrict our attention to the normally stable fast-slow system, which is a fast-slow system where the fast directions are stable.
[…]
A normally stable fast-slow system is a fast-slow system where the eigenvalues of the Jacobian, $D_{y}f(x,y,0)\in\mathbb{R}^{N_{y}\times N_{y}}$, are strictly to the left of the imaginary axis when $f(x,y,0)=0$.
[…]
(3) $M_{\epsilon}=\{(x,y):y=y^{*}(x,\epsilon)\},$
is locally invariant under (1). We define $M_{\epsilon}$ to be the slow manifold.
Here a bounded hypersurface is called a locally invariant manifold if the vector field defining the dynamical system is tangent to the hypersurface at all of its points [48]. This lemma is proved following Theorem 4 in [36]. The purpose of the FSNN is to be a universal approximator for normally stable fast-slow systems. The network is based on the following well known result due to Fenichel.
[…]
When $\epsilon>0$ is sufficiently small and $(x,y)$ lies is in a neighborhood around the slow manifold $M_{\epsilon}$, there exists a change of coordinates $(x,y)\rightarrow(\tilde{x},\tilde{y})$ where the dynamics of normally stable fast-slow systems can be written as,
(4a) $\frac{d}{dt}\tilde{y}$ $=A(\tilde{x},\tilde{y},\epsilon)\tilde{y},$
(4b) $\frac{d}{dt}\tilde{x}$ $=\epsilon{\tilde{g}}(\tilde{x},\tilde{y},\epsilon),$
where $A(\tilde{x},\tilde{y},\epsilon)\in\mathbb{R}^{N_{y}\times N_{y}}$, ${\tilde{g}}(\tilde{x},\tilde{y},\epsilon)\in\mathbb{R}^{N_{x}}$, and $A(\tilde{x},0,0)$ is a stable matrix, i.e., its eigenvalues are on the left of the imaginary axis.
The above lemma is proved in Chapter 3.2 of [36]. The lemma introduces (4), which is often called the Fenichel normal form due to the pioneering work of Neil Fenichel [3]. Note that the Fenichel normal form given in [36] is more general than (4), including fixed points that are either stable or unstable in the linearization of the fast variable. In this work, we only consider fast variables that are long-time stable. Moreover, due to the existence of the coordinate transformation, we consider the general form of the singularly perturbed systems (2). However, the advantage of the split form (1) is that the slow manifold dimension is known.
Our FSNN takes advantage of a modified formulation of Fenichel normal form which will be described in the following discussion. The following theorem forms the framework for the FSNN architecture.
[…]
Suppose the dynamics of $z(t;\epsilon)\in\mathbb{R}^{N}$, $\epsilon>0$, admit a fast-slow split and can be described by normally stable fast-slow system. Then there exists a $\epsilon$-dependent invertible transformation
(7) $\left[\begin{array}[]{c}y\\ x\end{array}\right]$ $=h(z;\epsilon),$
where $x(t)\in\mathbb{R}^{N_{x}}$, $y(t)\in\mathbb{R}^{N_{y}}$, are slow and fast variables, respectively, and for $|y|\leq\Delta$, where $\Delta>0$, the dynamics for $(x,y)$ can be described using
(8a) $\frac{d}{dt}y$ $=T(x)y+B(x,y,\epsilon)(y,y)+\epsilon C(x,y,\epsilon)y,$
(8b) $\frac{d}{dt}x$ $=\epsilon g(x,y,\epsilon).$
$T(x)\in\mathbb{R}^{N_{y}\times N_{y}}$ is a block upper triangular matrix consisting of either $\mathbb{R}^{2\times 2}$ or $\mathbb{R}$ blocks on the diagonal and the eigenvalues of these diagonal blocks are strictly to the left of the imaginary axis for all $x$. $B(x,y,\epsilon):\mathbb{R}^{N_{y}}\times\mathbb{R}^{N_{y}}\rightarrow\mathbb{R}^{N}_{y}$ is a bilinear map, $C(x,y,\epsilon)\in\mathbb{R}^{N_{y}\times N_{y}}$ is a matrix, and $g(x,y,\epsilon)\in\mathbb{R}^{N_{x}}$. $T$, $B$, $C,$ and $g$ are $C^{r}$ for $r\geq 1$.
In the following discussion, we will prove Theorem 8 by refining the Fenichel normal form in (4) and applying the matrix Schur decomposition. First, we modify (4) by performing a Taylor expansion of the term $A(\tilde{x},\tilde{y},\epsilon)\tilde{y}$ about $\tilde{y}=0$ and $\epsilon=0$. Assuming that $A$ is in $C^{r}$ for $r\geq 1$, we have
(9) $A(\tilde{x},\tilde{y},\epsilon)\tilde{y}=A(\tilde{x},0,0)\tilde{y}+B(\tilde{x},\tilde{y},\epsilon)(\tilde{y},\tilde{y})+\epsilon C(\tilde{x},\tilde{y},\epsilon)\tilde{y},$
where $B(\tilde{x},\tilde{y},\epsilon):\mathbb{R}^{N_{y}}\times\mathbb{R}^{N_{y}}\rightarrow\mathbb{R}^{N_{y}}$ is a bilinear map and $C(\tilde{x},\tilde{y},\epsilon)\in\mathbb{R}^{N_{y}\times N_{y}}$ are remainder terms. Next, the Schur decomposition of $A(\tilde{x},0,0)$ will be used to define a new set of fast coordinates.
[…]
The Schur form of a matrix, $A\in\mathbb{R}^{M\times M}$, is a matrix $T\in\mathbb{R}^{M\times M}$ that forms the Schur decomposition $A=QTQ^{T}$, where $Q\in\mathbb{R}^{M\times M}$ is orthogonal ($Q^{T}Q=I$) and $T$ is block upper triangular where the diagonal blocks consist of $\mathbb{R}$ and $\mathbb{R}^{2\times 2}$ blocks.
The existence of the Schur decomposition is proven in many references, including [49, 50], though for any given matrix, the decomposition is not unique. However, the eigenvalues of $A$ are given by the eigenvalues of the diagonal blocks of $T$. Let the Schur decomposition of $A(\tilde{x},0,0)$ be given by
(10) $A(\tilde{x},0,0)=Q(\tilde{x})T(\tilde{x})Q(\tilde{x})^{\rm T},$
where $Q(\tilde{x})\in\mathbb{R}^{N_{y}\times N_{y}}$ is an orthogonal matrix and $T(\tilde{x})\in\mathbb{R}^{N_{y}\times N_{y}}$ is the Schur form of $A(\tilde{x},0,0)$. Due to our assumptions on $A(\tilde{x},0,0)$, all of the eigenvalues of $T(\tilde{x})$ are to the left of the imaginary axis. Consider the coordinate transformation $\hat{y}=Q(\tilde{x})^{\rm T}\tilde{y}$. We form the relevant ODE for $\hat{y}$ to use in place of the ODE for $\tilde{y}$.
$\frac{d}{dt}\hat{y}=\;$ $Q(\tilde{x})^{\rm T}\frac{d}{dt}\tilde{y}+\left(\nabla Q(\tilde{x})^{\rm T}\right)\left(\frac{d}{dt}\tilde{x},\tilde{y}\right)$
=\;= ${Q(\tilde{x})^{T}}Q(\tilde{x})T(\tilde{x})Q(\tilde{x})^{T}\tilde{y}+Q(\tilde{x})^{\rm T}B(\tilde{x},Q(\tilde{x})\hat{y},\epsilon)(Q(\tilde{x})\hat{y},Q(\tilde{x})\hat{y})$
$+\epsilon Q(\tilde{x})^{\rm T}C(\tilde{x},\epsilon)Q(\tilde{x})\hat{y}+\epsilon\left(\nabla Q(\tilde{x})^{\rm T}\right)\left(g_{\epsilon}(\tilde{x},\hat{y}),\hat{y}\right).$
=\;= $T(\tilde{x})\hat{y}+Q(\tilde{x})^{\rm T}B(\tilde{x},Q(\tilde{x})\hat{y},\epsilon)(Q(\tilde{x})\hat{y},Q(\tilde{x})\hat{y})$
$+\epsilon Q(\tilde{x})^{\rm T}C(\tilde{x},\epsilon)Q(\tilde{x})\hat{y}+\epsilon\left(\nabla Q(\tilde{x})^{\rm T}\right)\left(g_{\epsilon}(\tilde{x},\hat{y}),\hat{y}\right).$
By redefining terms, it can be shown that the ODE is equivalent to
$\frac{d}{dt}\hat{y}=T(\tilde{x})\hat{y}+\tilde{B}(\tilde{x},\hat{y},\epsilon)(\hat{y},\hat{y})+\epsilon\tilde{C}(\tilde{x},\hat{y},\epsilon)\hat{y},$
where $\tilde{B}(\tilde{x},\hat{y},\epsilon)$ is a bilinear map and $\tilde{C}(\tilde{x},\hat{y},\epsilon)$ is a linear map. This establishes an equivalence to (8).
[…]
This theorem can be easily verified by substituting $y=0$ into (8a) and observing that $\frac{d}{dt}y=0$ on the slow manifold. Trajectories on the slow manifold will remain on the slow manifold for all time. In this case, the dynamics can be reduced to an ODE for the slow variable,
(11) $\frac{d}{d\tau}x=g(x,0,\epsilon),\qquad(x,y)\in M,$
[…]
(12) $d\left((x(t),y(t)),M\right)\leq\kappa\exp(\alpha t),$
where $d$ represents the Euclidean distance.
This theorem is proved following Theorem 5 in [36].
Theorem 8 provides a framework for the FSNN architecture, which expresses each of the functions $h,T,B,C,$ and $g$ using neural networks with specially designed architectures described in the following section. There are a few considerations that went into using the modified Fenichel normal form. By expanding $A(\tilde{x},\tilde{y},\epsilon)$ in (9), we expose three separate terms with known structures. We further use the Schur decomposition to explicitly expose the eigenvalues of $A(\tilde{x},0,0)$, which are required to be left of the imaginary axis. Finally, we further simplify the linear term by absorbing the matrix $Q(\tilde{x})$ into the representation. Though we use the Schur decomposition to represent $A(\tilde{x},0,0)$ in this paper, other decompositions for representing matrices with eigenvalues to the left of the imaginary axis have been studied in literature [51, 50].
3 Fast-Slow Neural Networks (FSNNs)
In this section, we present a neural network-based approach to approximating the set of solutions to normally stable fast-slow systems using the framework of Theorem 8. We refer to this approach as the fast-slow neural network (FSNN), which is defined below and summarized in Figure 1.
[…]
The fast slow neural network (FSNN) is a transformation of an initial condition, $z_{0}$, and parameter, $\epsilon>0$, to a time series $\{z_{n}\}_{n=0}^{N}$ corresponding to times $t_{0},\dots,t_{N}$, where $z_{n}\in\mathbb{R}^{N}$ for $n=0,\dots,N$. The initial condition is transformed into fast-slow coordinates, $(y_{0},x_{0})$, via (7) where $h$ is represented using an invertible coupling flow network. A numerical integration is used to approximate the solution to the fast-slow system in (8) to obtain the time series $\{(y_{n},x_{n})\}_{n=0}^{N_{t}}$ for corresponding times $t_{0},\dots,t_{N}$, where $T$ is represented using a negative Schur form network, $B$ is represented using a low-rank bilinear map network, and $C$ and $g$ are represented using feed-forward neural networks. The trajectories are then transformed back into the original coordinate system using
(14) $z_{n}=h^{-1}\left(\left[\begin{array}[]{cc}y_{n}&x_{n}\end{array}\right]^{\rm T};\epsilon\right),\qquad n=1,\dots,N.$
The key components of Definition 3, including the invertible coupling flow network, Schur form network, and low-rank bilinear map network, will be defined in the following sections. First we will begin by stating the main result of the paper.
[…]
The FSNN is a universal approximator for the set of solutions of normally stable fast-slow systems in the neighborhood of the slow manifold.
Due to Theorem 8, normally stable fast-slow systems can be represented using (7)–(8). Therefore, the FSNN inherits the universal approximation properties of its components, which we will later show are universal approximators for their corresponding class of functions. The invertible coupling flow network which represents $h$ is described in Section 3.1 and was proven to be a universal approximator to invertible maps in [37]. The universal approximation theorems of the negative Schur form network representing $T$ and the bilinear map network representing $B$ are proved in Sections 3.2 and 3.3, respectively. The universal approximation of nonlinear functions using feedforward neural networks for $C$ and $g$ are well known, e.g. [52]. Finally, we describe a consistent numerical scheme to obtain trajectories in time, which has an error bound that vanishes as the time steps $\Delta t_{n}=t_{n}-t_{n-1}\rightarrow 0$.
Our ML-based approach naturally comes with a set of hyperparameters. The standard hyperparmeters are those related to the architectures of $h$, $T$, $C$ and $g$. The only systemic hyperparameters are the dimensions of the slow and fast subspaces, $N_{x}$ and $N_{y}$, respectively. In the examples considered in Section 5, the true dimensions of slow manifolds are known and used. In applications when the size of the slow manifold dimension is not known, tools like diffusion maps [46] or proper orthogonal decomposition (POD) [53] may be used to infer the size of the reduced dimension before applying the proposed FSNN.
3.1 The Invertible Coupling Flow Network
Figure 1: Schematic of the fast-slow neural network (FSNN). The original coordinates are projected into fast-slow coordinates using the invertible coupling flow network. In these coordinates, numerical integration of the Fenichel normal form (8) is used to advance the solution to later time steps. The inverse function of the invertible coupling flow network is used to convert the trajectories in the fast-slow coordinates into the original coordinate system.
Now we will describe further details of the FSNN architecture. The first component is the invertible function, $h(z;\epsilon)$, which represents the transformation into fast-slow coordinates. The network is based on the invertible coupling flow network developed in [54], where the authors closely follow the architecture provided by [55]. This network is an example of an invertible neural network (INN). In particular, invertible coupling flow networks have been shown to be universal approximators to invertible functions in [37]. In this work, we modify the original network by introducing a regularization property which controls the Lipschitz constant of both the forward and inverse map. Such a regularization improves the robustness of the network during training. The constant is a tunable parameter of the network that, similar to the hidden dimensions of the network, can be made arbitrarily large to increase the expressiveness of the network. Therefore, this modification is compatible with the universal approximation result derived in [37]. As such, we introduce the key ingredient of Lipschitz control.
[…]
The bi-Lipschitz affine transformation (bLAT) layer is given by the transformation, $f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{M}$, defined by
(15) $f(x)=U\Sigma V^{\rm T}x+b,$
where $U\in\mathbb{R}^{M\times r}$, $V\in\mathbb{R}^{N\times r}$, $\Sigma\in\mathbb{R}^{r\times r}$, $b\in\mathbb{R}^{M}$, $r=\min(M,N)$, $U$ and $V$ are orthogonal matrices satisfying $U^{\rm T}U=V^{\rm T}V=I$, and
(17) $\Sigma={\rm diag}\left[\begin{array}[]{ccc}\sigma_{1}&\dots&\sigma_{r}\end{array}\right],$
where $\frac{1}{L}\leq\sigma_{i}\leq L$ for $i=1,\dots,r$ and some $L\geq 1$.
We remark that when $M=N$, the mapping is invertible.
(18) $f^{-1}(x)=V\Sigma^{-1}U^{\rm T}(x-b).$
Our new contribution is to include the matrix term $U\Sigma V^{T}$, which is written as a singular value decomposition (SVD), which exists for every matrix in $\mathbb{R}^{M\times N}$. One advantage of the bLAT layer is the following theorem.
[…]
The forward and inverse bLAT layers are both $L$-bi-Lipschitz.
This follows from the fact that $\frac{1}{L}\leq\|\nabla f(x)\|_{2}\leq L$ and $\frac{1}{L}\leq\|\nabla f^{-1}(x)\|_{2}\leq L$. Another advantage of using the SVD is that, by construction, the evaluation of $f(x)$ and $f^{-1}(x)$ have an equivalent computational cost. In our implementation of the $M=N$ case, the orthogonal matrices of the bLAT layers are parameterized using the matrix exponential of a skew-symmetric matrix input, i.e. $U={\rm expm}(S)$, where $S=-S^{T}$. When $M\neq N$, the orthogonal matrices are parameterized using the Householder factorization.
The second component of the invertible coupling flow network is the affine coupling flow layer. We adapt the definition provided in Section 2.1 of [37] and introduce a similar regularization idea to the above.
[…]
The affine coupling flow layer is defined by the transformation $y={{g}}(x)$, $x,y\in\mathbb{R}^{M}$, which is implicitly defined by
(27) $\left[\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right]=\left[\begin{array}[]{cc}{\rm diag}(F(x_{2}))&\\ &{\rm diag}(G(y_{1}))\end{array}\right]\left[\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right]+\left[\begin{array}[]{c}B(x_{2})\\ C(y_{1})\end{array}\right],$
where $x=\left[\begin{array}[]{cc}x_{1}&x_{2}\end{array}\right]^{\rm T}$ $y=\left[\begin{array}[]{cc}y_{1}&y_{2}\end{array}\right]^{\rm T}$ $x_{1},y_{1}\in\mathbb{R}^{d}$, $x_{2},y_{2}\in\mathbb{R}^{M-d}$, $F,B:\mathbb{R}^{M-d}\rightarrow\mathbb{R}^{d}$, $G,C:\mathbb{R}^{d}\rightarrow\mathbb{R}^{M-d}$, $1\leq d<M$ and the range of the output elements of $F$ and $G$ is constrained to $\left[\frac{1}{L},\,L\right]$ for some $L\geq 1$. The inverse mapping, $x=f^{-1}(y)$, is implicitly defined by
(36) $\left[\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right]=\left[\begin{array}[]{cc}{\rm diag}(F(x_{2}))^{-1}&\\ &{\rm diag}(G(y_{1}))^{-1}\end{array}\right]\left(\left[\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right]-\left[\begin{array}[]{c}B(x_{2})\\ C(y_{1})\end{array}\right]\right).$
The functions $F,G,B,C$ are represented using feed-forward neural networks. In order to constrain outputs to the range $\left[\frac{1}{L},L\right]$, we apply the map $\theta_{L}:\mathbb{R}\rightarrow\left[\frac{1}{L},L\right]$ defined by $\theta_{L}(s)=\exp(\log(L)\tanh(s))$. This constraint ensures invertibility of the affine coupling layer. The Lipschitz constant is a function of $L$ and $\nabla F,\nabla G,\nabla B,\nabla C$ and can optionally be constrained using bLAT layers in the feedforward networks.
The invertible coupling flow network combines bLAT layers and affine coupling flow layers into a network.
[…]
Let $f_{1},\dots f_{K}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ define a set of bLAT layers and $g_{1},\dots g_{K}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ define a set of affine coupling flow layers. The invertible coupling flow network, $h:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ is defined as the composition
(37) $h=g_{K}\circ f_{K}\circ\dots\circ g_{1}\circ f_{1}$
and the inverse map is defined by
(38) $h^{-1}=f_{1}^{-1}\circ g_{1}^{-1}\circ\dots\circ f_{K}^{-1}\circ g_{K}^{-1}.$
The free parameter $L$ in each of the layers adjusts the complexity of the overall network and can be used as an effective regularizer. Despite the introduction of this regularization parameter, the network can be made arbitrarily expressive with the tuning of $L$, and therefore the universal approximation result from [37] still holds true.
[…]
The invertible coupling flow network is a universal approximator for invertible maps.
In Theorem 8, the invertible transformation is $\epsilon$-dependent. Therefore, we incorporate this dependence into the network by including $\epsilon$ as a functional dependence in $F,G,B,C$.
3.2 The Negative Schur Form Network
Next, a parameterization suitable for representing $T$ is discussed. We present the negative Schur form network, which is a differentiable parameterization for Schur forms with eigenvalues to the left of the imaginary axis. The Schur form network is composed of a feedforward neural network in which the outputs are re-formatted into a negative Schur form matrix, which is defined below.
[…]
The negative Schur form matrix is given by the block matrix $T\in\mathbb{R}^{N_{y}\times N_{y}}$, where
(43) $T=\left[\begin{array}[]{cccc}T_{11}&T_{12}&\dots&T_{1p}\\ &T_{22}&\dots&T_{2p}\\ &&\ddots&\vdots\\ &&&T_{pp}\end{array}\right].$
When $M$ is even, $p=M/2$ and $T_{ij}\in\mathbb{R}^{2\times 2}$ for $i=1,\dots p$, $j=i,\dots,p$. When $M$ is odd, $p=(M+1)/2$, $T_{ij}\in\mathbb{R}^{2\times 2}$ for $i=1,\dots p-1$, $j=i,\dots,p-1$, $T_{ip}\in\mathbb{R}^{2}$ for $i=1,\dots,p-1$, and $T_{pp}\in(-\infty,0)$. Each of the $\mathbb{R}^{2\times 2}$ blocks on the diagonal are further parameterized using the map
(48) $\mathcal{M}(R,r,\theta,\phi)=R\left[\begin{array}[]{cc}\cos(\theta)&\sin(\theta)\\ -\sin(\theta)&\cos(\theta)\end{array}\right]+r\left[\begin{array}[]{cc}\cos(\phi)&\sin(\phi)\\ \sin(\phi)&-\cos(\phi)\end{array}\right],$
where the region of parameters is restricted to
(49) $|R|\geq|r|,\quad R<0,\quad\theta\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right),\quad{{\phi\in\mathbb{R}}}$
Due to this parameterization, we have the following theorem.
[…]
The negative Schur form network is a parameterization for the set of Schur forms with eigenvalues to the left of the imaginary axis.
This can be proven by showing that (48)–(49) can smoothly represent both complex conjugate pairs and real pairs of eigenvalues with negative real part. Equation (48) represents the $\mathbb{R}^{2\times 2}$ matrix as a 4-parameter family sum of a symmetric and antisymmetric operator [56]. Let $m=\frac{1}{2}{\rm Tr}(\mathcal{M})=R\cos(\theta)$ and $p={\rm det}(\mathcal{M})=R^{2}-r^{2}$. Then the eigenvalues of $\mathcal{M}$ are given by
(50) $\lambda_{\pm}=m\pm\sqrt{m^{2}-p}.$
The real parts of the eigenvalues are given by
(51) ${\rm Re}(\lambda_{\pm})=m\pm\sqrt{\max(m^{2}-p,0)}.$
Figure 2 shows the regions of the $m{\rm-}p$ plane where both eigenvalues are less than zero (green region), greater than zero (red region), complex (blue region), and real (outside of the blue region). In order for ${\rm Re}(\lambda_{\pm})<0$, we must have $m<0$ and $p>0$. In order to enforce these constraints, we enforce the restrictions in (49). We remark that for the first constraint, $R\cos(\theta)<0$, $R$ and $\cos(\theta)$ must have opposite signs. If we focus on the region $\theta\in(-\pi/2,\pi/2)$ where $\cos(\theta)>0$, we would require $R<0$. If the region of $\theta$ were instead shifted by $\pi$, then we would require $R<0$. However, since $\mathcal{M}(R,r,\theta+\pi,\phi)=\mathcal{M}(-R,r,\theta,\phi)$, these choices are equivalent.
Figure 2: Plot showing the eigenvalues encountered in various regions of the $m$-$p$ plane, where $m=\frac{1}{2}{\rm Tr}(\mathcal{M})=R\cos(\theta)$ and $p={\rm det}(\mathcal{M})=R^{2}-r^{2}$. The green region corresponds to linear stability, ${\rm Re}(\lambda_{+},\lambda_{-})<0$, and the red region corresponds to linear instability, ${\rm Re}(\lambda_{+},\lambda_{-})>0$. Above the parabola $p=m^{2}$ (in the blue region), eigenvalues are complex. Below the parabola (outside of the blue region), the eigenvalues are real.
In the implementation of the network, each of the $\mathbb{R}^{2\times 2}$ blocks have four tunable parameters, $\bar{R},\,\bar{r},\,\bar{\theta},\,\bar{\phi}\in\mathbb{R}$, that are mapped onto the restricted domain using
(52) $R=-|\bar{r}|e^{|\bar{R}|},\quad r=\bar{r},\quad\theta=\frac{\pi}{2}\tanh(\bar{\theta}),\quad\phi=\bar{\phi}.$
The remaining off-diagonal terms are unrestricted in parameter space. In order to achieve dependence with respect to $x$, (i.e. $T(x)$), we introduce a feedforward network which takes $x$ as an input and outputs both the tunable parameters $\bar{R},\,\bar{r},\,\bar{\theta},\,\bar{\phi}$ for each diagonal block and the elements of the remaining blocks. The resulting negative Schur form matrix is returned as the output of $T(x)$.
3.3 The Low-Rank bilinear map Network
Finally, we introduce an approximation for the bilinear map term, $B$.
[…]
The low-rank bilinear form network $B:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{y}}\rightarrow\mathbb{R}^{N_{z}}$ is defined by
(53) $B(x,y)=\sum_{r=1}^{R}(C^{(r)}x)\odot(D^{(r)}y),$
where $x\in\mathbb{R}^{N_{x}}$, $y\in\mathbb{R}^{N_{y}}$, $C^{(r)}\in\mathbb{R}^{N_{z}\times N_{x}}$, $D^{(r)}\in\mathbb{R}^{N_{z}\times N_{y}}$, $\odot$ represents the Hadamard product (element-wise multiplication), and $R$, $1\leq R\leq\min(N_{x},N_{y})$, represents the rank of approximation.
Here, $R$ is a tunable parameter. Increasing $R$ gives more expressive power at the expense of computational cost. This network has the following property:
[…]
The low-rank bilinear map network is a universal approximator to the set of bilinear forms.
Consider the bilinear form, $\tilde{B}:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{y}}\rightarrow\mathbb{R}^{N_{z}}$. The tensor representation of $\tilde{B}$ is written as $\tilde{B}_{ijk}$ where $\tilde{B}_{i}\in\mathbb{R}^{N_{x}\times N_{y}}$ for $i=1,\dots,N_{z}$. Let $p=\min(N_{x},N_{y})$, then we can use the singular value decomposition to represent the matrix $\tilde{B}_{i}$ as
(54) $\tilde{B}_{i}=\sum_{r=1}^{p}\sigma_{i}^{(r)}u_{i}^{(r)}(v_{i}^{(r)})^{T},$
where $\sigma_{i}^{(r)}$, $u^{(r)}$, $v^{(r)}$, $r=1,\dots,p$, are the singular values, left singular vectors, and right singular vectors of $\tilde{B}_{i}$, respectively, and $\sigma_{1}\geq\dots\geq\sigma_{p}\geq 0$. The $i^{\rm th}$ term of $\tilde{B}(x,y)$ is given by
(55) $x^{T}\tilde{B}_{i}y$ $=\sum_{r=1}^{p}\sigma_{i}^{(r)}x^{T}u_{i}^{(r)}(v_{i}^{(r)})^{T}y,=\sum_{r=1}^{p}\sigma_{i}^{(r)}\left(\sum_{j=1}^{N_{x}}u_{ij}^{(r)}x_{j}\right)\left(\sum_{k=1}^{N_{y}}v_{ik}^{(r)}y_{k}\right),$
We can find some $C\in\mathbb{R}^{N_{z}\times N_{x}}$ and $D\in\mathbb{R}^{N_{z}\times N_{y}}$ such that
(56) $x^{T}\tilde{B}_{i}y$ $=\sum_{r=1}^{p}\left(\sum_{j=1}^{N_{x}}C_{ij}^{(r)}x_{j}\right)\left(\sum_{k=1}^{N_{y}}D_{ik}^{(r)}y_{k}\right).$
This is equivalent to
(57) $\tilde{B}(x,y)$ $=\sum_{r=1}^{p}(C^{(r)}x)\odot(D^{(r)}y).$
The network uses a truncated version of this sum in equation (53) to approximate the action of the bilinear map $\tilde{B}$. The approximation error of the low-rank bilinear map network is $\|B_{i}-\tilde{B}_{i}\|_{2}=\sigma_{i}^{(R+1)}$. This error can be decreased by increasing $R$.
In our formulation, the operator $B$ has coefficients which are functions of $x,y,$ and $\epsilon$. We inject this functional dependence by making the elements of $C^{(r)}$ and $D^{(r)}$ outputs of a feed-forward network with inputs $(x,y,\epsilon)$.
3.4 Numerical Integration
The Schur form, $T$ will have eigenvalues that can be of arbitrarily large magnitude. Due to this, the dynamics of (8) will be stiff. Therefore, an Implicit-Explicit (IMEX) time integrator can be used to solve (8) efficiently. The IMEX scheme has been applied to a neural ODE framework in [57]. As shown below, an additional benefit of the proposed FSNN is to eliminate the need for iterative or direct linear solvers, thereby simplifying the training process.
Consider an IMEX Runge-Kutta method consisting of a diagonally implicit method for treating the stiff linear term and an explicit method for treating the nonlinear term. The method can be represented by the following Butcher tableau [58]
$a_{11}$
$a_{21}$ $a_{22}$ $\alpha_{21}$
$\vdots$ $\vdots$ $\ddots$ $\vdots$ $\ddots$
$a_{s1}$dots $\dots$ $a_{ss}$ $\alpha_{s1}$dots $\alpha_{s,{s-1}}$
$b_{1}$dots $\dots$ $b_{s}$ $\beta_{1}$dots $\beta_{s-1}$ $\beta_{s}$
Let $x_{n}\approx x(t_{n})$, $y_{n}\approx y(t_{n})$, and $\Delta t=t_{n}-t_{n-1}$. Let $f(x,y,\epsilon)=B(x,y,\epsilon)(y,y)+\epsilon C(x,y,\epsilon)y$. The intermediate stages are implicitly defined by
(58a) $X^{i}$ $=x_{n-1}+\epsilon\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{g(X^{j},Y^{j},\epsilon)},$
(58b) $Y^{i}$ $=y_{n-1}+\Delta t\sum_{j=1}^{i}a_{ij}{T(X^{j})Y^{j}}+\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{f(X^{j},Y^{j},\epsilon)}$
for $i=1,\dots,s$. In our notation, we take the above summation when $i=1$ to be 0 (i.e., $\sum_{j=1}^{0}(\cdot)=0$). Equation (58b) can be rewritten to solve for $Y^{i}$.
(59) $Y^{i}$ $=(I-\Delta ta_{ii}T(X^{i}))^{-1}\left(y_{n-1}+\Delta t\sum_{j=1}^{i-1}\left(a_{ij}{T(X^{j})Y^{j}}+\alpha_{ij}{f(X^{j},Y^{j},\epsilon)}\right)\right)$
Note that the matrix given by $I-\Delta ta_{ii}T(X^{i})$ is a block 2-by-2 upper triangular matrix. Due to the proposed architecture, this inverse can be handled efficiently using a backward substitution solve. Since $T$ additionally has functional dependence, we treat this in a vectorization-friendly matrix-free fashion, which avoids storing multiple matrices $T(X^{i})$ for each data pair in the batch set during training. The solution at the next time step is obtained through the update
(60a) $x_{n}$ $=x_{n-1}+\epsilon\Delta t\sum_{i=1}^{s}\beta_{i}{g(X^{i},Y^{i},\epsilon)},$
(60b) $y_{n}$ $=y_{n-1}+\Delta t\sum_{i=1}^{s}(b_{i}{T(X^{i})Y^{i}}+\beta_{i}{f(X^{i},Y^{i},\epsilon)}).$
In our implementation, we use the two-stage second-order accurate IMEX strong-stability-preserving L-stable scheme defined in [58].
When the solution is initially on the slow manifold ($y=0$), the stages $Y^{i}$ for $i=1,\dots,s$ evaluate exactly to zero and therefore the numerical scheme can be simplified to an integration of the slow variables given by
(61a) $x_{n}$ $=x_{n-1}+\epsilon\Delta t\sum_{i=1}^{s}\beta_{i}{g(X^{i},0,\epsilon)},$
(61b) $X^{i}$ $=x_{n-1}+\epsilon\Delta t\sum_{j=1}^{i-1}\alpha_{ij}{g(X^{j},0,\epsilon)},\qquad i=1,\dots,s.$
In this case, one can take larger time steps (e.g. $\Delta t\sim\frac{1}{\epsilon}$) and step over the fast dynamics formally. A hybrid scheme can also be used where the full system in (60) can be integrated until $\|y\|$ is sufficiently small, after which slow integration in (61) is pursued using larger time steps. We note that some of the first works to use Runge-Kutta based residual networks to learn equations from time series was pursued in [59, 60].
Data-driven closure
A closure discovery procedure is implied through the above integration with large time steps. For a given full state vector $z_{0}$, the transformation $h$ first maps to the separate state of $(x_{0},y_{0})$. The state vector is then projected onto the closure of $y=0$; the system for $x$ is thus closed and evolved using (61). Finally, the evolved state of $x$ is lifted back to the full state $z$ using $h^{-1}$.
4 Numerical Implementation
Figure 3: Schematic of training the fast-slow neural network (FSNN). The network is trained on dynamics data, $\mathcal{D}$, which contains examples of trajectories on both the fast and slow timescales and manifold data, $\mathcal{M}$, which contains examples of solution vectors that lie on the slow manifold. This data interacts both with the full network and the component involving the invertible neural network (INN). The overall loss is a combination of the dynamics loss, which includes $L_{2}$ losses in the original coordinates ($\mathcal{L}_{\rm system}$), fast coordinates ($\mathcal{L}_{\rm fast}$), and slow coordinates ($\mathcal{L}_{\rm slow}$), and the manifold loss ($\mathcal{L}_{\rm manifold}$), which is an $L_{2}$ loss of the fast coordinate, which should map to $y=0$ on the slow manifold.
In this section, we describe the numerical implementation of the fast-slow neural network and its corresponding training algorithm. The computation of the loss function and its gradient is summarized in Figure 3. The training set consists of both trajectory data, which describes the behavior of the dynamics, and manifold data, which describes the limiting behavior of the system. The trajectory dataset, $\mathcal{D}$, contains time series, $\{(t_{n}^{(k)},z_{n}^{(k)})\}_{n=1}^{N_{t}}$, and parameters $\epsilon^{(k)}$, for $k=1,\dots,N_{d}$, where $N_{d}$ is the number of trajectory data samples. The manifold dataset, $\mathcal{M}$, contains multiple examples of the solution on the slow manifold, $z^{(k)}$, with corresponding parameters $\epsilon^{(k)}$, for $k=1,\dots,N_{m}$. Examples of manifold data can be obtained either experimentally by integrating a model for a sufficiently long time with a small $\epsilon$ or analytically if the equations permit an $\epsilon=0$ reduction or asymptotic expansion. For many examples using the zeroth order expansion with $\epsilon=0$ is sufficient. Often obtaining a higher order expansion is nontrivial.
The loss function is given by
(62) $\mathcal{L}=\mathcal{L}_{\rm system}+\mathcal{L}_{\rm fast}+\mathcal{L}_{\rm slow}+\mathcal{L}_{\rm manifold}.$
The first three terms are evaluated using the trajectory data set, $\mathcal{D}$, and the last term is evaluated using the manifold data set, $\mathcal{M}$. For the dynamics-oriented loss terms, the fast-slow neural network is applied to initial conditions in $\mathcal{D}$ to obtain the model trajectories, $\{(t_{n}^{(k)},\tilde{x}_{n}^{(k)},\tilde{y}_{n}^{(k)},\tilde{z}_{n}^{(k)})\}$. Additionally, the invertible coupling flow network is applied to the trajectory in $\mathcal{D}$ to augment the data $\{(t_{n}^{(k)},{x}_{n}^{(k)},{y}_{n}^{(k)},{z}_{n}^{(k)})\}$. Let $L$ be an operator representing the approximation to the time integral of squared error, given by
(63) $L(t,z,\tilde{z})=\sum_{n=0}^{N_{t}-1}\frac{1}{2}\left((z_{n+1}-\tilde{z}_{n+1})^{2}+(z_{n}-\tilde{z}_{n})^{2}\right)(t_{n+1}-t_{n}).$
Then we choose
(64a) $\mathcal{L}_{\rm system}$ $=\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}L(t,z^{(k)},\tilde{z}^{(k)}),$
(64b) $\mathcal{L}_{\rm fast}$ $=\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}L(t,y^{(k)},\tilde{y}^{(k)}),$
(64c) $\mathcal{L}_{\rm slow}$ $={{\frac{1}{N_{d}}\sum_{k=1}^{N_{d}}\frac{L(t,x^{(k)},\tilde{x}^{(k)})}{\epsilon^{(k)}}.}}$
Including the three terms individually encourages the network to learn the dynamics for the original and fast-slow coordinates to an adequate level.
In practice, we split the trajectory dataset, $\mathcal{D}$, into a fast-scale subset, with time steps chosen to smoothly capture the fast dynamics converging to the slow manifold, and a slow-scale subset, with larger time steps chosen to capture the variation of the slow variables after the dynamics sufficiently close to the slow manifold. The trajectories in fast-scale subset are treated using the full integration in (60) while the trajectories from the slow-scale subset are treated using the slow-manifold integration in (61).
For the manifold loss term, the invertible coupling flow network is applied to examples in $\mathcal{M}$ to obtain fast-slow coordinates $(x^{(k)},y^{(k)})$. If the data is on the slow manifold, then $y^{(k)}$ should be zero. We therefore include the following as a penalty term.
(65) $\mathcal{L}_{\rm manifold}=\frac{1}{N_{m}}\sum_{k=1}^{N_{m}}(y^{(k)})^{2}$
Our implementation of the fast-slow neural network uses Jax [61] as a backend. Auto-differentiation is used to compute gradients of the loss function with respect to the tunable parameters and the Adam optimizer from the Optax [62] package is used adjust the weights towards an locally optimal solution for the dataset.
5 Examples
In this section, we consider multiple example problems to demonstrate the applicability of the FSNN to various problems, including multi-scale problems arising in hydrodynamics and plasma physics. In training, we explored various combinations of meta parameters for each problem. We pursued a policy to train several models in parallel and choose the model with the best performance. The meta parameters for the best models are reported in each respective section. We remark that though increasing the number of network parameters increases the expressiveness of the model, it comes at a higher computational cost and increases the risk of overfitting.
5.1 Visualization of the Attracting Slow Manifold Property
A significant advantage of the FSNN is that, for any choice of model parameters, the network will have an attracting slow manifold that, due to Theorem 8, will attract solutions for sufficiently small values of $\epsilon$. The slow manifold is represented by the graph, $y=0$. In Figure 4, we demonstrate this property on multiple random initializations of FSNNs for $N_{y}=N_{x}=1$. In this figure, network a, b, and c are representing different FSNNs with different weights and therefore different slow manifolds. Multiple initial conditions are drawn randomly from a standard normal distribution (left plot) and are integrated using each FSNN to a later time $t=4$. The trajectory points at $t=4$ (blue dots) and the resulting slow manifold (orange curves) are plotted at this final time, demonstrating convergence of solutions to the slow manifolds inherent to each network.
Figure 4: Demonstration of attracting slow manifold property of the FSNN. For three different random initializations of the FSNN, initial conditions are drawn randomly from a standard normal distribution and are evolved in time to a later time ($t=4$). The dynamics are shown to limit to the curve representing the slow manifold.
5.2 Learning a Simple Attracting Manifold
We demonstrate the ability of the FSNN to learn dynamics for a simple test problem where the dynamics can be described by the following system of ODEs for $z_{1},z_{2}\in\mathbb{R}$.
(66a) $\frac{dz_{1}}{dt}$ $=-\epsilon(\sin(z_{1})+z_{2})$
(66b) $\frac{dz_{2}}{dt}$ $=\lambda(z_{1})(z_{2}-\theta(z_{1}))$
In this example, $z_{1}$ represents the slow variable and $z_{2}$ represents the fast variable. When $\epsilon=0$, the slow manifold is given by $\theta(z_{1})$. In this example, we choose $\lambda$ and $\theta$ to be
(67) $\lambda(x)=-1-\frac{1}{10}\cos(2x),\qquad\theta(x)=2\tanh(x).$
The trajectory dataset, $\mathcal{D}$, consists of $10^{3}$ fast-scale trajectories, corresponding to initial conditions for $z_{1},z_{2}$ randomly generated uniformly in $[-1,1]$, $\epsilon$ randomly generated log-uniformly in $[10^{-5},10^{-2}]$, and integrated forward 5 timesteps using an adaptive RK45 solver with an error tolerance of $10^{-8}$. The manifold dataset, $\mathcal{M}$, consists of $10^{3}$ examples of data pairs $((z_{1},z_{2}),\epsilon=0)$ on the graph $z_{2}=\theta(z_{1})$, where $z_{1}$ is randomly generated uniformly in $[-1,1]$.
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 10. Each of the feedforward neural networks on the RHS of the neural ODE system consists of 10 layers and the bilinear map has a rank of 2.
Figure 5 shows a comparison of the trained model and the reference solution in phase space for various choices of $\epsilon$. As can be seen by the corresponding error plots, the model shows reasonable agreement to the reference solution. In all of the phase plots, it is evident that the trajectories eventually get trapped in a slow manifold. In the rightmost plot on the top row, a comparison of the learned eigenvalues with that of the ground truth, $\lambda(z_{1})$, is shown. Additionally in the rightmost plot on the bottom row, the error between the learned slow manifold and true slow manifold is shown. The results demonstrate that the FSNN is able to successfully recover details of the original operator.
Figure 5: Results of learning the dynamics of the simple attracting manifold example described by (66). Left three plots on top row: phase portraits comparing the ground truth (g.t.) reference solution (bold curves) and that of the trained model (dots) for $\epsilon=0,0.001,0.01$. Left three plots on bottom row: corresponding errors in time for each trajectory in the above phase portraits. Rightmost plot on top row: comparison of the eigenvalue of the negative Schur form as a function of $z_{1}$ learned by the FSNN with the ground truth eigenvalue given by $\lambda$ in (67). Rightmost plot on bottom row: error between the learned slow manifold and the ground truth slow manifold given by $\theta$ in (67).
5.3 Linear Grad Moment System
In this example, we show that the FSNN can be applied towards learning hydrodynamic slow manifolds. We consider a linear Grad moment system [63, 33], defined by
(68a) $\partial_{t}\rho$ $=-\epsilon\,\nabla\cdot\mathbf{u},$
(68b) $\partial_{t}\mathbf{u}$ $=-\epsilon\left(\nabla\rho+\nabla T+\nabla\cdot\sigma\right),$
(68c) $\partial_{t}T$ $=-\frac{2\epsilon}{3}\left(\nabla\cdot\mathbf{u}+\nabla\cdot\mathbf{q}\right),$
(68d) $\partial_{t}\mathbf{\sigma}$ $=-\mathbf{\sigma}-\epsilon\left(2\overline{\nabla\mathbf{u}}+\frac{4}{5}\overline{\nabla\mathbf{q}}\right),$
(68e) $\partial_{t}\mathbf{q}$ $=-\frac{2}{3}\mathbf{q}-\epsilon\left(\frac{5}{2}\nabla T+\mathbf{\nabla}\cdot\mathbf{\sigma}\right),$
where $\rho(\mathbf{x},t)$ is the density, $\mathbf{u}(\mathbf{x},t)$ is velocity, $T(\mathbf{x},t)$ is temperature, $\sigma(\mathbf{x},t)$ is the stress tensor, $\mathbf{q}(\mathbf{x},t)$ is the heat flux, and $\mathbf{x}=(x,y,z)^{T}\in\mathbb{R}^{3}$ is the spatial coordinate. The overline represents the symmetric, traceless part of a tensor, e.g.
(69) $\overline{\mathbf{a}}$ $=\frac{1}{2}(\mathbf{a}+\mathbf{a}^{T})-\frac{1}{3}\mathbf{I}\,\text{Tr}(\mathbf{a}).$
We consider solutions that are one-dimensional ($\partial_{y}=\partial_{z}=0$) and $2\pi$-periodic in space, e.g.
(70) $\rho(\mathbf{x},t)=\sum_{k=-\infty}^{\infty}\rho_{k}(t)e^{ikx}.$
The Fourier coefficients are governed by the following system of ODEs.
(86) $\frac{d}{dt}\left[\begin{array}[]{c}\rho_{k}\\ u_{k}\\ T_{k}\\ \sigma_{k}\\ q_{k}\end{array}\right]=\left[\begin{array}[]{ccccc}0&-ik\epsilon&0&0&0\\ -ik\epsilon&0&-ik\epsilon&-ik\epsilon&0\\ 0&-\frac{2}{3}ik\epsilon&0&0&-\frac{2}{3}ik\epsilon\\ 0&-\frac{4}{3}ik\epsilon&0&-1&-\frac{8}{15}ik\epsilon\\ 0&0&-\frac{5}{2}ik\epsilon&-ik\epsilon&-\frac{2}{3}\end{array}\right]\left[\begin{array}[]{c}\rho_{k}\\ u_{k}\\ T_{k}\\ \sigma_{k}\\ q_{k}\end{array}\right],\quad k=0,\pm 1,\dots$
For this example, we truncate the Fourier series to consider only the $k=0,\pm 1$ models. Initial conditions are randomly generated for each Fourier coefficient uniformly in the complex region $[-1,1]\times[-i,i]$. The solution is computed for 50 time steps both using $\Delta t_{\rm fast}=\frac{1}{4}$ and $\Delta t_{\rm slow}=\frac{1}{4\epsilon}$. The model is trained using only the information from the first time step and tested using the remaining of the trajectory. The slow-scale trajectories are offset $\Delta t_{\rm slow}$ from $t=0$ to ensure the solution is sufficiently close to the slow manifold. A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 40. The feedforward neural networks on the RHS of the neural ODE system each consist of 50 layers and the bilinear map has a rank of 1. Considering each coefficient and its real and imaginary parts, the dimension of the learned model is $\mathbb{R}^{30}$, with $N_{x}=18$ and $N_{y}=12$.
Table 1: Comparison of ground truth eigenvalues of the Jacobian of the fast equations with that of the negative Schur form of the trained model for the Grad Moment System problem.
Grad Eigenvalues
g.t. model error g.t. model error
-1 -0.9971 2.9e-03 -$\frac{2}{3}$0.6659 8.2e-04
-1 -0.9969 3.1e-03 -$\frac{2}{3}$0.6658 8.3e-04
-1 -0.9969 3.1e-03 -$\frac{2}{3}$0.6658 9.1e-04
-1 -0.9968 3.2e-03 -$\frac{2}{3}$0.6657 9.3e-04
-1 -0.9968 3.2e-03 -$\frac{2}{3}$0.6657 9.4e-04
-1 -0.9968 3.2e-03 -$\frac{2}{3}$0.6657 9.4e-04
Figure 6 shows comparisons of the trained model and reference solutions on the fast time scale and slow time scale. The results demonstrate that reasonable accuracy can be achieved on both time scales. We observe that model for the slow dynamics is slightly off in phase which causes error to grow in time.
Table 1 shows a comparison of the eigenvalues of the trained model with the analytical eigenvalues. Quantitative agreement is observed, which implies that the model was successful in mapping the slow manifold approximately to the true slow manifold.
It is critical to stress that the slow dynamics comparison in Figure 6 is done through integration on the slow manifold using (61). Specifically, the initial condition state vector $z_{0}$ is first transformed into the fast-slow coordinates $(x_{0},y_{0})$ and projected onto the learned slow manifold $(x_{0},0)$ before integrating the slow variable $x$ on the slow time scale only and then transforming back into original coordinates using $h^{-1}$. This approach is significantly more efficient than integrating the full system using a standard numerical integrator. The excellent accuracy of the predicted slow variable solution indicates a data-driven closure is learned through this procedure. The closure of (68) formally exists when $\epsilon$ approaches 0. It is readily seen that the reduced system is a linearized Euler system. When $\epsilon=0$, the critical manifold is simply the center subspace. When weak nonlinearity is added, there are center manifolds tangent to the center subspace. In general these center manifolds are not unique, but there is a unique smoothest center manifold known as the spectral submanifold, see [64, 65]. In future work, it would be interesting to explore whether our methods can be adapted to ensure the learned slow manifold agrees with the spectral submanifold.
Figure 6: Comparison of fast scale (top) and slow scale (bottom) dynamics between the ground truth and trained model for the Grad moment system example.
5.4 Multi-Scale Lorenz96 Equations
In this section we consider the two-scale Lorenz96 equations, which was a model proposed in [66] as a test bed model for multi-scale behavior in atmospheric dynamics. Multiscale techniques for this system have been explored in many works, such as in [67]. The original system is formulated as
(87a) $\frac{d}{d\tau}X_{k}$ $=-X_{k-1}(X_{k-2}-X_{k+1})-X_{k}+F-(hc/b)\sum_{j=1}^{J}Y_{j,k},$
(87b) $\frac{d}{d\tau}Y_{j,k}$ $=-cbY_{j+1,k}(Y_{j+2,k}-Y_{j-1,k})-cY_{j,k}+(hc/b)X_{k},$
for $j=1,\dots,J$ and $k=1,\dots,K$, where the solutions satisfy periodicity ($X_{k+K}=X_{k}$, $Y_{j,k+K}=Y_{j,k}$, $Y_{j+J,k}=Y_{j,k}$). Let $c=b=1/\epsilon$, $h=1$, $Y=\epsilon y$, $\tau=\epsilon t$, and $F=0$. The system becomes
(88a) $\frac{d}{dt}x_{k}$ $=-\epsilon x_{k-1}(x_{k-2}-x_{k+1})-\epsilon x_{k}-\epsilon^{2}\sum_{j=1}^{J}y_{j,k},$
(88b) $\frac{d}{dt}y_{j,k}$ $=-y_{j+1,k}(y_{j+2,k}-y_{j-1,k})-y_{j,k}+x_{k}.$
Since the transformation is singular at $\epsilon=0$, we study the equations in the region $0<\epsilon\ll 1$.
[…]
Using Definition 2, we need to show that the Jacobian of the right-hand-side of (88b) has eigenvalues strictly to the left of the imaginary axis. Consider the following nonlinear equation $f(y;\xi)\in\mathbb{R}^{J}$ for $y\in\mathbb{R}^{J}$ and $\chi\in\mathbb{R}$.
(89) $f(y;\chi)=-y_{j+1}(y_{j+2}-y_{j-1})-y_{j}+\chi.$
The Jacobian of this equation, $D_{y}f$ is given by
(90) $(D_{y}f)_{jk}=\frac{\partial f_{j}}{\partial y_{k}}=-\delta_{jk}-\delta_{{j+1},k}(y_{j+2}-y_{j-1})-y_{j+1}(\delta_{j+2,k}-\delta_{j-1,k}),$
where $\delta_{jk}$ represents the Kronecker delta. We are interested in the eigenvalues of $D_{y}$ when $f(y;\chi)=0$. Note that one solution to $f(y;\chi)=0$ is given by $y_{j}=\chi$ for $j=1,\dots,J$. At this fixed point, the Jacobian becomes
(91) $(D_{y}f)_{jk}|_{y=\xi}=-\delta_{jk}-\chi(\delta_{j+2,k}-\delta_{j-1,k})$
This represents a circulant matrix, of which the eigenvalues are given by
(92) $\lambda_{j}=-1-\chi(e^{ij4\pi/J}-e^{-ij2\pi/J}),\qquad j=0,1,\dots,J-1.$
Note that
(93) ${\rm Re}(\lambda_{j})=-1-\chi(\cos(4\xi_{j})-\cos(2\xi_{j})),$
where $\xi_{j}=4\pi j/J$. We are interested in the region for $\chi$ where ${\rm Re}(\lambda_{j})<0$ for all $j$. As $J\rightarrow\infty$, we have the bound $-\frac{9}{8}\leq\cos(4\xi_{j})-\cos(2\xi_{j})\leq 2$. This implies that the region is given by $-\frac{1}{2}<x<\frac{8}{9}$. For a finite $J$, the bounds of $\cos(4\xi_{j})-\cos(2\xi_{j})$ are determined by the resolution of $\xi_{j}$, which is a discrete function of $j$. This affords us to take a slightly wider region of $x$ where ${\rm Re}(\lambda_{j})<0$ for all $j$. For example, when $J=4$, $-\frac{1}{2}<x<1$.
Homotopy continuation can be used on the polynomial system of equations defined by $f$ to reveal a high-order polynomial containing all possible roots. Maple was used to evaluate all the roots for multiple values of $\chi$ and for $J=4,5,6$. Within the region $\chi\in(-\frac{1}{2},\frac{8}{9})$, the only real-valued solution is given by $y_{j}=\chi$ for $j=1,\dots,J$. We conjecture that there is only a single real-valued solution to (89) when $\chi\in(-\frac{1}{2},\frac{8}{9})$ for $J>6$.
We have shown that the multi-scale Lorenz96 equations are normally-stable for a restricted range of $x_{k}$. However, for any $x_{k}$, $y_{j,l}$, $\epsilon>0$, the system is energy stable, which implies that, for any initial condition, the energy will decay in time and the dynamics will eventually be able to be described by a normally fast-slow system for some time $t>0$.
[…]
The equations defined by (88) are energy-stable for $\epsilon>0$.
Consider again the original system. We multiply the equations by $x_{k}$ and $y_{j,k}$, respectively. Let $(\cdot)^{\prime}=\frac{d}{dt}(\cdot)$, then
(94a) $x_{k}x_{k}^{\prime}$ $=-\epsilon x_{k}x_{k-1}(x_{k-2}-x_{k+1})-\epsilon x_{k}^{2}-\epsilon^{2}\sum_{j=1}^{J}x_{k}y_{j,k}$
(94b) $y_{j,k}y_{j,k}^{\prime}$ $=-y_{j,k}y_{j+1,k}(y_{j+2,k}-y_{j-1,k})-y_{j,k}^{2}+x_{k}y_{j,k}.$
We collect the terms and take the sum over the discrete variables.
$\frac{1}{2}\sum_{k=1}^{K}(x_{k}^{2})^{\prime}$ $=-\epsilon\sum_{k=1}^{K}x_{k}x_{k-1}x_{k-2}+\epsilon\sum_{k=1}^{K}x_{k+1}x_{k}x_{k-1}-\epsilon\sum_{k=1}^{K}x_{k}^{2}-\epsilon^{2}\sum_{k=1}^{K}\sum_{j=1}^{J}x_{k}y_{j,k}$
$\frac{1}{2}\sum_{j=1}^{J}(y_{j,k}^{2})^{\prime}$ $=-\sum_{j=1}^{J}y_{j,k}y_{j+1,k}y_{j+2,k}+\sum_{j=1}^{J}y_{j-1,k}y_{j,k}y_{j+1,k}-\sum_{j=1}^{J}y_{j,k}^{2}+\sum_{j=1}^{J}x_{k}y_{j,k}$
Due to periodicity, the cubic terms evaluate to zero. The remaining terms give us the following energy estimate.
(95) $\frac{1}{2}\sum_{k=1}^{K}(x_{k}^{2})^{\prime}+\frac{1}{2}\epsilon^{2}\sum_{k=1}^{K}\sum_{j=1}^{J}(y_{j,k}^{2})^{\prime}=-\epsilon\sum_{k=1}^{K}x_{k}^{2}-\epsilon^{2}\sum_{k=1}^{K}\sum_{j=1}^{J}y_{j,k}^{2}<0$
We train a FSNN to learn the multi-scale Lorenz96 equations from data choosing $J=K=4$. The trajectory dataset, $\mathcal{D}$, consists of $10^{4}$ fast-scale trajectories, corresponding to initial conditions for $x_{k}$ randomly generated uniformly in $[-\frac{1}{2},1]$ and $y_{j,k}$ randomly generated from a normal distribution, $\epsilon$ randomly generated log-uniformly in $[10^{-5},10^{-2}]$, and integrated forward 10 timesteps using an adaptive RK45 solver with error tolerance $10^{-8}$. The manifold dataset, $\mathcal{M}$ consists of $10^{4}$ examples of data pairs $((x_{k},y_{j,k}),\epsilon=0)$ on the graph $y_{j,k}=x_{k}$, where $x_{k}$ is randomly generated uniformly in $[-\frac{1}{2},1]$.
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 40. The feedforward neural networks on the RHS of the neural ODE system each consist of 40 layers and the bilinear map has a rank of 4.
Figure 7: The top five plots are trajectory plots comparing the reference solution computed using RK45 (bold curves) and that of the trained model (dotted markers) for an initial condition with $\epsilon=1.1\cdot 10^{-3}$ for the multi-scale Lorenz96 example described by the dynamics in (88). The bottom five plots show the corresponding errors as a function of time. The plots in the left column show the solution and error on the slow time scale ($t\in[0,3000]$) while the plots in the right column zoom in on the fast time scale ($t\in[0,3]$).
Figure 7 shows a comparison of the trained model and the reference solution for an example initial condition with $\epsilon=1.1\cdot 10^{-3}$. The model shows quantitative agreement to the reference solution. Although the globally attracting solution is the zero solution, the error remains away from 0. This can be explained by approximation error of the model and can be improved through including more data in the training set. Figure 8 shows a comparison of the eigenvalues of the trained model with the analytical eigenvalues. Agreement is observed, implying that the model was successful in mapping the slow manifold approximately to the true slow manifold. In this example, the time interval of all the training data is much shorter than $t=3$. Nevertheless, the prediction can be done up to $T=3000$ or longer with excellent accuracy.
Figure 8: Comparison of the ground truth eigenvalues of the Jacobian of the fast dynamics (bold lines) with that of the negative Schur form learned from the trained model (dots) for the multi-scale Lorenz96 example described by the dynamics in (88)
5.5 Abraham-Lorentz Equations
In this section, we consider the Abraham-Lorentz equations [1], which model the radiation reaction in the case of a single electron with charge $e$ and mass $m$ moving through a static external magnetic field $\mathbf{B}(\mathbf{x})\in\mathbb{R}^{3}$. The equations are given by
(96a) $\frac{2}{3}\frac{e^{2}}{c^{3}}\frac{d\mathbf{a}}{d\tau}$ $=m\mathbf{a}-\frac{e}{c}\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(96b) $\frac{d\mathbf{v}}{d\tau}$ $=\mathbf{a},$
(96c) $\frac{d\mathbf{x}}{d\tau}$ $=\mathbf{v},$
where $\mathbf{x}(\tau),\mathbf{v}(\tau),\mathbf{a}(\tau)\in\mathbb{R}^{3}$ are the position, velocity, and acceleration, respectively, and $c$ is the speed of light. We rescale these equations following the approach in [1]. Time is scaled by the observer time scale $T$ as $\tau=T\overline{\tau}$, space by the observer length scale $\mathbf{x}=L\overline{\mathbf{x}}$, velocity as $\mathbf{v}=(L/T)\overline{\mathbf{v}}$, the magnetic field as $\mathbf{B}(\mathbf{x})=B_{0}\overline{\mathbf{B}}(\mathbf{x})$, and acceleration as $a=(L/T)(|e|B_{0})(mc)^{-1}\overline{\mathbf{a}}.$ We reformulate the rescaled Abraham-Lorentz equations in terms of the nondimensional variables and drop the overline for notational convenience.
(97a) $\left(\frac{r_{0}}{cT}\right)\frac{2}{3}\frac{d\mathbf{a}}{d\tau}$ $=\mathbf{a}-\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(97b) $\frac{d\mathbf{v}}{d\tau}$ $=(\omega_{c}T)\mathbf{a},$
(97c) $\frac{d\mathbf{x}}{d\tau}$ $=\mathbf{v},$
where we have introduced the cyclotron frequency $\omega_{c}=|e|B_{0}/(mc)$, the classical electron radius $r_{0}=e^{2}/(mc^{2})$, and the sign of the charge $\zeta\in\{-1,+1\}$, where $\zeta=-1$ corresponds to electrons and $\zeta=+1$ to positrons. The dimensionless parameters
(98) $\gamma_{R}=\frac{r_{0}}{cT},\qquad\gamma_{B}=\frac{1}{\omega_{c}T},$
represent the ratio of the electron size to the distance light travels during a time interval $T$, and ratio of the cyclotron period to $T$, respectively. The system can be transformed into a two-scale fast-slow system with the choice $\gamma_{B}=1$ and $\gamma_{R}~{}=~{}\epsilon~{}\ll~{}1$. We scale time through the transformation $\tau=\frac{2}{3}\epsilon t$ to reveal a normally-unstable fast-slow system given by
(99a) $\frac{d\mathbf{a}}{dt}$ $=\mathbf{a}-\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(99b) $\frac{d\mathbf{v}}{dt}$ $=\epsilon\frac{2}{3}\mathbf{a},$
(99c) $\frac{d\mathbf{x}}{dt}$ $=\epsilon\frac{2}{3}\mathbf{v},$
We will refer to (99) as the forward-time Abraham-Lorentz system. A normally-stable fast-slow system can be obtained by reversing time using $t_{-}=-t$,
(100a) $\frac{d\mathbf{a}}{dt_{-}}$ $=-\mathbf{a}+\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x}),$
(100b) $\frac{d\mathbf{v}}{dt_{-}}$ $=-\epsilon\frac{2}{3}\mathbf{a},$
(100c) $\frac{d\mathbf{x}}{dt_{-}}$ $=-\epsilon\frac{2}{3}\mathbf{v}.$
We refer to (100) as the reverse-time Abraham-Lorentz system. In both systems, $\mathbf{a}$ is the fast variable and $\mathbf{v},\mathbf{x}$ are the slow variables. In reverse time, the acceleration, $\mathbf{a}$ will be attracted to the slow manifold, given by
(101) $\mathbf{a}^{*}(\mathbf{x},\mathbf{v})=\zeta\mathbf{v}\times\mathbf{B}(\mathbf{x})+\epsilon\frac{3}{2}\mathbf{v}\times\mathbf{B}(\mathbf{x})\times\mathbf{B}(\mathbf{x})+\mathcal{O}(\epsilon^{2}).$
Though the reverse-time system is stable on the fast time scale, the system exhibits an instability tangential to the slow manifold for which solutions blow up on the slow time scale. This is a more challenging test than the previous ones.
It is of interest to obtain forward-time trajectories of electrons along the slow manifold. A standard numerical integration of (99) will result in unstable trajectories on the fast time scale since any numerical error will push trajectories off of the slow-manifold, resulting in exponential growth of the acceleration. We instead apply FSNNs to learn the Fenichel normal form of the system in reverse time. We then obtain forward-time trajectories along the learned slow manifold by performing an integration of the slow variables along $y=0$ using (61). This test is therefore a more practical application of a data-driven closure.
We train an FSNN to learn the Abraham-Lorentz equations from data. The trajectory dataset, $\mathcal{D}$, consists of $10^{4}$ trajectories, corresponding to initial conditions randomly generated in the region, $\mathcal{R}$, specified by
=$\mathcal{R}=\left\{\mathbf{x},\mathbf{v},\mathbf{a}:\quad 2\leq r\leq 4,\quad|z|\leq 1,\quad\|\mathbf{v}\|\leq 1,\quad\|\mathbf{a}-\mathbf{a}^{*}(\mathbf{x},\mathbf{v})\|\leq 1\right\},$
where $r=\sqrt{x_{1}^{2}+x_{2}^{2}}$ and $z=x_{3}$. $\epsilon$ is randomly generated log-uniformly in the range $[10^{-5},10^{-2}]$. The trajectories are integrated in reverse time using equations (100) using an adaptive RK45 solver set to an error tolerance of $10^{-8}$. The first 40 time steps batched in sequences of length 10 are used as fast-scale training data. These trajectories are also interpolated over a slower time scale with $\Delta t=0.025/\epsilon$ and $10^{4}$ sequences of length 10 are chosen at random for slow-scale training data. The manifold dataset, $\mathcal{M}$ consists of $10^{4}$ examples of data pairs $((\mathbf{x},\mathbf{v},\mathbf{a}),\epsilon=0)$ on the graph $\mathbf{a}=\mathbf{v}\times\mathbf{B}(\mathbf{x})$, where $\mathbf{x},\mathbf{v}$ are randomly generated in the region $\mathcal{R}$.
The magnetic field is chosen to be $\mathbf{B}=B_{r}\mathbf{e}_{r}+B_{\phi}\mathbf{e}_{\phi}+B_{z}\mathbf{e}_{z}$, where $\mathbf{e}_{r},\mathbf{e}_{\phi},\mathbf{e}_{z}$ represent basis vectors for cylindrical coordinates, and
(102) $B_{r}$ $=-\frac{z}{q(r,z)r}B_{0},\qquad B_{\phi}=\frac{R_{0}}{r}B_{0},\qquad B_{z}=\frac{(r-R_{0})}{q(r,z)r}B_{0},$
where $q(r,z)=q_{0}+q_{2}((r-R_{0})(r-R_{0})+z^{2})/a^{2}$ and $q_{0}=1.2$, $q_{2}=2.8$, $a=1.5$, $R_{0}=3.0$, $Z_{0}=3.0$, and $B_{0}=1.0$. This field is an idealized equilibrium of a tokamak, where the field is physical inside a torus with major radius of $R_{0}$ and minor radius of $a$. Here $q$ is called a safety factor [68] and the poloidal magnetic flux function is $\psi=\frac{B_{0}a^{2}}{2q_{2}}\ln(q).$ Under this definition, the given field can be rewritten as
(103) $\mathbf{B}=\frac{1}{R}\nabla\psi\times\mathbf{e}_{\phi}+\frac{F_{0}}{R}\mathbf{e}_{\phi}.$
where the poloidal field is a constant of $F_{0}=R_{0}B_{0}$. This test thus models a charged particle moving through a tokamak with radiation reaction.
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 50. The function $g$ on the right-hand-side of the neural ODE uses a single-layer feedforward neural network with hidden dimension 100 and is added to a bilinear form network with rank 1. The remaining feedforward neural networks on the right-hand-side of the neural ODE system each consist of 50 layers and the bilinear map has a rank of 4.
Figure 9 shows comparisons of the solutions in $r$-$z$ space between the trained network and trajectories computed using RK45 on the reverse-time system. Using the reverse-time trajectories, endpoints were supplied as initial conditions to the trained network to predict forward in time. For both the reverse-time predictions (red arrows) and forward-time predictions (green arrows), the trained model agrees well with the solution computed numerically. The corresponding $L_{2}$ errors are shown for position, velocity, and acceleration in Figure 10, demonstrating that errors are low when integrating on the slow time scale.
Table 2: Comparison of the true eigenvalues Jacobian of the fast dynamics to the eigenvalues of the negative Schur form of the trained model for the Abraham-Lorentz problem.
Reverse-Time Abraham-Lorentz System Eigenvalues
g.t. model error
-1 -0.99962 3.8e-04
-1 -0.99958 4.2e-04
-1 -0.99958 4.2e-04
Table 2 shows the comparison of the true eigenvalues of the linear part of the fast dynamics of the reverse-time system. The eigenvalues were recovered with a great deal of accuracy, demonstrating that the fast-scale dynamics are learned accurately.
Figure 11 shows comparisons of orbits between the trained model and orbits obtained when integrating the system by substituting the first two terms of the asymptotic expansion in (101). Note again that when integrating along the slow manifold, only the slow variable is evolved using the transformation $h$ and the closure $y=0$. The trained model is able to qualitatively capture the gyro-motion behavior while also accurately capturing the guiding center dynamics over the slow time scale. Despite training using trajectories over a relatively short time frame (2 or less gyration periods) due to the slow time scale instability of the reverse-time system, the trained network is highly capable of predicting long-time dynamics in forward-time for many more gyration cycles. We remind the reader that standard numerical methods cannot integrate the forward-time Abraham-Lorentz system along the slow manifold without experiencing an exponential growth instability. Direct integration of the asymptotic expansion removes the instability, but introduces an error dependent on $\epsilon$ due to the asymptotic approximation. This result highlights the unique capability of the proposed network.
Figure 9: Displayed are the trajectories for three values of $\epsilon$ evolved backwards in time along the slow manifold (top three plots, red arrows) and three values of $\epsilon$ evolved forwards in time along the slow manifold (bottom three plots, green arrows). The solid lines represent the ground truth trajectories and the dotted lines represent the model.
Figure 10: Error trajectories for the reverse time (top three plots) and forward time simulations are displayed. For each simulation, the $L_{2}$ norm of position, velocity and acceleration are plotted.
Figure 11: Comparison of orbits of the forward-time Abraham-Lorentz system (99) along the slow manifold for the first two terms of the asymptotic expansion in (101) (left) and the trained model (right) for two different initial conditions and choices of $\epsilon$.
6 Conclusions
In this paper, we introduced the fast-slow neural network (FSNN), which is a data-driven approach suitable for learning singularly perturbed dynamical systems where the fast-scale dynamics are dissipative. Our method enforces the existence of a trainable, attracting invariant slow manifold as a hard constraint. Invertible neural networks and neural ODEs are key components of the FSNN. We introduce bi-Lipschitz affine transformation (bLAT) layers for enforcing regularity of the invertible neural network. We also develop a negative Schur form network which parameterizes matrices with eigenvalues to the left of the imaginary axis. This in addition to the use of an additive L-stable diagonally-implicit Runge Kutta scheme for integration are key pieces in ensuring that the fast-scale dynamics are stable. The Schur form network also enables the use of back-substitution to efficiently handle the implicit solve. Analytical representation of the slow manifold enables efficient integration on the slow time scale. We demonstrate the FSNN on many examples that exhibit two timescales, including the Grad moment system from hydrodynamics, two-scale Lorenz96 equations for modeling atmospheric dynamics, and Abraham-Lorentz dynamics modeling radiation reaction of electrons in a magnetic field.
In order to effectively train the FSNN, a sufficient number of trajectories across multiple values of $\epsilon$ and initial conditions are necessary. Despite this, we demonstrate that the network can be used to discover a full-order model with long-term accuracy near a slow manifold when only short duration trajectories are available. Therefore the proposed network is also applicable to discover a reduced-order model as a unique closure discovery procedure.
In the examples we consider, the equations and slow manifold reduction in the limit $\epsilon=0$ are known a priori. Therefore, we were able to bypass an important shortcoming of our approach, which is the need to determine the dimensionality of the slow manifold. While a plethora of techniques exist in the literature [47], it would be interesting to combine them to our technique to tackle problems where the dimensionality is unknown.
A future direction of our work involves considering hyperbolic fast slow systems, which can be represented using a more general Fenichel normal form consisting of both stable and unstable directions.
Acknowledgment
This research used resources provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001, and the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0023112.